# Autonomous Horizon-based Asteroid Navigation With Observability-constrained Maneuvers

### v1

@@ 6.2.1 Monte Carlo Analysis
The monte carlo analysis is repeated for the approach and circularization scenario. The same settings from Section 6.1.1 are used, with initial position standard deviation of $\sigma=30$ meters and 200 iterations. The regular controller succeeds in 68 out of 200 iterations for a success rate of 34%. The EKF and true dynamics diverge often and lead to the spacecraft either crashing into the asteroid or traveling out of the system. The observability-constrained Lyapunov controller shows improved performance with a success rate of 181 out of 200 iterations for a success rate of 90.5%. Meanwhile, the same case with a [2.5 1 1] ellipsoid succeeds for 168 out of 200 iterations for a success rate of 84%. This supports the consistent performance benefit of the observability-constrained controller across multiple scenarios. The controller is shown to be useful in approach scenario at larger distance, where the ellipsoid assumption better holds. While the performance degrades for the ellipsoid test in this case as well, it maintains better performance than the regular controller. All Monte Carlo results are shown in Fig. 16.
Figure 16: Monte Carlo Simulation of 200 True Trajectories using Basic Lyapunov controller (Left), Observability-constrained Lyapunov controller with Spherical (Middle) and [2.5 1 1] Ellipsoid Primary (Right). Initial position (green circle), final position (red cross) and orbital direction (black arrows) indicated.
@@ 7 Discussion
This paper presents a promising approach to autonomous asteroid OpNav using a Lyapunov controller with path constraints for observability-constrained maneuvers. The approach offers a combination of OpNav methods, state estimation, and control, especially when navigating challenging environments around asteroids.
@@
The developed controller overcomes the limited applicability of the Horizon-based OpNav algorithm under extreme lighting conditions for asteroid navigation. When the spacecraft is positioned behind the asteroid’s dark side, there is poor observability for the system due to an undetectable horizon, potentially leading to inaccurate state estimations. The observability-constrained controller concept allows the design of a control scheme that ensures a spacecraft follows a trajectory that avoids these faulty measurements. This method is better than simply discarding poor measurements, which may lead to an extended gap without measurements, causing the state estimate to diverge.
This research also develops an understanding of the applicability of horizon-based OpNav as a measurement for asteroid navigation. It is found that it is effective up to a range of 25-30 times a target asteroid’s radius, outside of a cone with a 30-degree half-angle from either straight-on or dark-side viewing, and displays reasonable performance with an ellipsoidal target body up to a [5,1,1] shape axis ratio. While the image generation method used for this research is quick, generating each image at an average rate of 0.3 seconds is the primary computational restriction for the Monte Carlo simulation.
@@
While effective under certain conditions, the implementation shows a divergence in scenarios with extreme noise, highly nonlinear dynamics, or erroneous initial states. This highlights either a possible need for further tuning and refinement of the EKF or a limitation due to the controller not accounting for navigation errors. However, it is effective within realistic bounds and expectations of these parameters that would be found in an actual space mission. The system demonstrates effective state estimation for both spherical and ellipsoidal bodies while the spacecraft is in stable orbits. However, the performance deteriorates with more complex, nonlinear trajectories or highly elliptical asteroids.
This solution does not yet provide sufficient observability for shape estimation, particularly when attempting to estimate the radius or shape ratios of the asteroid. More information is required for the filter to estimate these parameters successfully. Two options are through additional optical methods to infer properties such as the asteroid’s triaxial ellipsoid shape ratios and differential imaging to determine velocity to make the system observable in terms of shape estimation. These methods are a current area of investigation and the next step in this algorithm’s development.
@@
This method of using a horizon-based OpNav with observability-constrained Lyapunov control may offer a more computationally efficient and autonomous option for the early mission approach phase. It reduces the need for human-in-the-loop OpNav since the spacecraft can navigate itself to attain useful measurements as it approaches the asteroid. Furthermore, it has computation efficiency and accuracy advantages over other OpNav methods for a wider range of observable distances and angles. However, it is important to note that the proposition to use horizon-based OpNav with observability-constrained maneuvers is not intended to replace the methodologies used in existing space missions fully. Instead, we offer an additional method that is feasible to implement to improve mission efficiency and safety. This is because the observability constraint ensures continuous and reliable measurements at all mission stages while within the horizon-based OpNav range. The Lyapunov control and state estimation algorithms combined with the OpNav enable autonomy across the early mission approach to close proximity phases of an asteroid mission. There is potential for future work to extend its applicability, improve performance in extreme conditions, and develop more sophisticated state and shape estimation methods. These improvements are critical for expanding the system’s use in future deep space exploration missions.
@@ 8 Conclusion
This research successfully demonstrates a robust autonomous navigation and control method for spacecraft approaching or conducting operations near asteroids. A comprehensive simulation environment is developed that considers spacecraft dynamics around an asteroid, synthetic asteroid imaging and processing, horizon-based OpNav, EKF, and observability-constrained Lyapunov control. The novel contribution is a Lyapunov controller with path constraints to improve horizon-based OpNav measurement observability. Its effectiveness has been demonstrated by testing its stability when combined with EKF-based state estimation within acceptable error bounds. The validation with a simplified model of Bennu as a benchmark highlights the reliable test of its capabilities in this application. The system’s robustness in varying scenarios, including near-linear hyperbolic trajectories, flybys, and various orbit transfers, highlights its potential for broader applications in autonomous asteroid exploration.
Appendix A Path Constraint Partial Derivatives
@@
This section extends the derivation in Section 4.1 by computing the partial derivatives of the path constraints. The path constraint partial derivatives in terms of Milankovitch orbital elements are as follows:
@@
Appendix B Controller Gain Tuning Guidelines
Since the gains determined for the controller used in this research are for specific test cases with Bennu, it is imperative to develop guidelines to derive it for different scenarios to make it useful for diverse missions. After investigating the relations between the gains and the controller performance, the following observations and guidelines are made:
@@ Appendix B Controller Gain Tuning Guidelines
This Lyapunov controller formulation may be adapted for different use cases using these observations. Results demonstrating the effectiveness of these gains can be seen in Section 6.

### v2

@@ 5.2.1 Monte Carlo Analysis
The monte carlo analysis is repeated for the approach and circularization scenario. The same settings from Section˜5.1.1 are used, with initial position standard deviation of $\sigma=30$ meters and 200 iterations. The regular controller succeeds in 68 out of 200 iterations for a success rate of 34%. The EKF and true dynamics diverge often and lead to the spacecraft either crashing into the small body or traveling out of the system. The observability-constrained Lyapunov controller shows improved performance with a success rate of 181 out of 200 iterations for a success rate of 90.5%. Meanwhile, the same case with a [2.5 1 1] ellipsoid succeeds for 168 out of 200 iterations for a success rate of 84%. This supports the consistent performance benefit of the observability-constrained controller across multiple scenarios. The controller is shown to be useful in approach scenario at larger distance, where the ellipsoid assumption better holds. While the performance degrades for the ellipsoid test in this case as well, it maintains better performance than the regular controller. All approach and circularization Monte Carlo results are shown in Fig.˜12.
Similarly as the previous scenario, the evolution of EKF position Norm Error, OpNav measurement Norm Error and angle constraint satisfaction over time across the unconstrained, constrained with sphere and constrained with ellipse test cases are displayed in Fig.˜13. All of the trajectories in the unconstrained case violate the angle constraint, which corresponds to a significantly higher measurement error than the constrained cases with poorly characterized covariance. Thus, this leads to significant divergence in the unconstrained case. Both constrained cases offer significant improvements, with the results for the spherical body being better than the ellipsoid body due to lower error magnitude. However, the EKF covariance poorly bounds the errors in the basic controller case, corresponding to a deviation between the truth and EKF position estimate, which is concordant with our earlier definition of divergence. Furthermore, the constraints are correctly satisfied by the observability-maintaining Lyapunov controller as the unconstrained scenario enters the dark side cone with an angle lower than $30^{\circ}$ while the constrained cases remain consistently above. For the unconstrained case, the Lyapunov function value starts at a low value before increasing due to the natural progression of the trajectory, after which the controller attempts to reach the target orbit, in the process of which it encounters divergence since it does not consider the angle constraint. Meanwhile, the constrained cases start with a higher initial value allowing the controller to apply initial input to avoid violating the angle constraint, which then leads to two peaks (relatively delayed compared to the unconstrained peaks) that ensure further angle constraint satisfaction. This allows the controller to maintain trajectories that maintain observability, thereby minimizing divergence. This is concordant with the continuous thrust profiles depicting behaviour with an initial control peak, following by final circularization and angle constraint satisfaction inputs. The unconstrained trajectory encounters the maximum thrust constraint due to divergence and therefore has a significantly higher $\Delta v$ usage, while the constrained trajectories have a reasonable thrust distribution.
(a) Unconstrained
(b) Constrained (Spherical primary)
(c) Constrained ([2.5 1 1] Ellipsoid primary)
Figure 12: Approach and Circularization scenario. Monte Carlo simulation of 200 true trajectories using (a) basic Lyapunov controller, (b) observability-constrained Lyapunov controller with spherical primary, and (c) with [2.5 1 1] ellipsoid primary. Initial position (green circle), final position (red cross), and orbital direction (black arrows) are indicated.
Figure 13: Approach and Circularization scenario. Monte Carlo Simulation of EKF Position Norm Error (Row 1), OpNav Measurement Norm Error (Row 2) and Angle Constraint Satisfaction (Row 3) with 200 Trials using Basic Lyapunov Controller (Column 1), Observability-constrained Lyapunov Controller with Spherical (Column 2) and [2.5 1 1] Ellipsoid Primary (Column 3).
@@ 6 Discussion
The numerical results demonstrate that incorporating observability path constraints into a Lyapunov controller can improve autonomous asteroid OpNav performance compared to an unconstrained baseline. The developed controller maintains OpNav observability while considering the perturbation sensitivities of trajectories in low-gravity small-body environments. When the spacecraft is positioned behind the asteroid’s dark side, there is poor observability for the system due to an undetectable horizon, potentially leading to inaccurate state estimations. The observability-constrained controller concept allows the design of a control scheme that ensures a spacecraft follows a trajectory that avoids these faulty measurements. This method is better than simply discarding poor measurements, which may lead to an extended gap without measurements, causing the state estimate to diverge.
In both orbit-maintenance and approach–circularization scenarios, the constrained controller consistently maintained spacecraft trajectories within the observability and safety region, avoiding cone, range or collision violations observed in the unconstrained case. This stability was preserved across spherical and rotating ellipsoidal bodies, even under significant initial dispersions and geometry variations.
Across both testing scenarios, orbit maintenance and approach & circularization, the constrained controller satisfied the keep-out cone and range constraints throughout successful runs, while the unconstrained Lyapunov baseline frequently violated the angle constraint and diverged. In the stationkeeping case, the constrained controller achieved a success rate of 98% over 200 Monte Carlo trials with a spherical primary (vs. 4% for the unconstrained baseline), and 80.5% for a rotating ellipsoid. In the approach case starting near the range limit, the constrained controller achieved 90.5% success for the sphere (vs. 34% baseline) and 84% for the ellipsoid. These outcomes align with the time histories: (i) the angle trajectories for the constrained runs remained above the $30^{\circ}$ half-angle threshold, (ii) the EKF position errors were predominantly bounded by the $3\sqrt{\lambda_{\max}(P_{r})}$ envelopes, and (iii) the Lyapunov function exhibited characteristic peaks coincident with constraint activation, followed by monotone decrease as the trajectory re-entered the interior of the safe set. The controller maintained performance under modeling and operational variations such as rotating ellipsoids inducing periodic geometry variations, high initial state dispersion, and challenging target orbits that are near the observability constraint.
It is important to note that this controller formulation is for continuous, low-thrust control. As determined in Section˜5, the maximum instantaneous acceleration across both scenarios was $6.6203\times 10^{-6}\,\textrm{m/s}^{2}$, which aligns with low-thrust capabilities [74]. The observability constraint also aids the controller in maintaining a thrust profile below the spacecraft’s capabilities, while the unconstrained case leads to frequent thrust maximizations. Furthermore, the observability-constrained controller maintains consistently low $\Delta v$ levels within spacecraft capabilities.
@@
While effective under certain conditions, the implementation shows a divergence in scenarios with extreme noise, highly nonlinear dynamics, or erroneous initial states. This highlights either a possible need for further tuning and refinement of the EKF or a limitation due to the controller not accounting for navigation errors. However, it is effective within realistic bounds and expectations of these parameters that would be found in an actual space mission. The system demonstrates effective state estimation for both spherical and ellipsoidal bodies while the spacecraft is in stable orbits.
@@ 7 Conclusion
This research successfully demonstrates a robust autonomous navigation and control method to maintain OpNav observability for spacecraft approaching or conducting operations near small bodies. A comprehensive simulation environment is developed that considers spacecraft dynamics around an asteroid, synthetic asteroid imaging and processing, horizon-based OpNav, EKF, and observability-constrained Lyapunov control. The novel contribution is a Lyapunov controller with path constraints to improve horizon-based OpNav measurement observability. Its effectiveness has been demonstrated by testing controller stability when combined with EKF-based state estimation within acceptable error bounds.
We conduct extensive empirical validation demonstrating onboard feasibility. In two representative proximity scenarios, the controller delivered substantial gains in success rate over an unconstrained Lyapunov baseline. It improved from 4% to 98% in stationkeeping and 34% to 90.5% in approach for a spherical primary while preserving constraint satisfaction throughout successful runs and producing low-thrust, low-$\Delta v$ profiles compatible with electric propulsion. Comparable performance for rotating ellipsoids (80.5% and 84% success) supports robustness to target geometry and operational variability. The analysis incorporates Monte Carlo statistics, Lyapunov function evolution, position error bounds, thrust histories and $\Delta v$ distribution histograms, providing a complete picture of stability, constraint adherence, and control effort across diverse mission conditions. The validation with simplified sphere and ellipse small-body models as a benchmark highlights the reliable test of its capabilities in this application. The system’s robustness in varying scenarios highlights its potential for broader applications in autonomous small body exploration.
Appendix A Synthetic Image Generation Pipeline
The image generation is done using built-in MATLAB graphics tools. The procedure involved illuminating a white sphere or ellipsoid against a black background using an infinitely far away light source that is aligned to originate from the negative $x$-axis in the Hill frame to represent a parallel ray light source like the Sun. The camera is then placed based on an input distance, azimuth, and elevation calculated from the current spacecraft position and an image is captured. Gaussian white noise is added to the image to add realistic complexity for the filter to process. Edge detection is executed by identifying high gradient magnitudes between pixels on the image, which is commonly where the lit limb edge meets blank space. The generated images were compared with those produced by existing higher fidelity image generators in identical conditions and found to be sufficiently accurate when comparing the resulting OpNav measurement.
For the simulations in this paper, we generate the spheres and ellipsoids using the mean radius of any sample small body. Below is a stepwise summary of the process utilized for the synthetic image generation:
  1. 1.
    Generate a sphere or ellipsoid scaled according to the camera settings and distance or an experimentally determined scale factor.
  2. 2.
    Rotate the object using the small body’s axis, rotation period and simulation time elapsed. This information is used to calculate and apply the angle by which the object had rotated at each time step during simulation with the equation $\theta_{\textrm{sb}}=\frac{t}{t_{\textrm{rp}}}360^{\circ}$ where $\theta_{\textrm{sb}}$ is the angle by which the small body has rotated, $t$ is the time elapsed and $t_{\textrm{rp}}$ is the small body’s synodic rotation period.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
    Set the figure size and axis limits according to the image size, apply the camera’s FOV setting to the axes’ camera view angle, and set aspect ratios to be equal.
  7. 7.
  8. 8.
    Add white gaussian noise if desired. For this study, $\sigma=0.01$ of normalized noise was applied to the images.
The synthetic image generation process is also expressed in Algorithm˜1:
Algorithm 1 Image Generation Procedure
1:procedure imgimg=ImageGenerator(x_y_z,up,dist,FOV,imgSize,rad,ellip_shape,R_ellipx\_y\_z,up,dist,FOV,imgSize,rad,ellip\_shape,R\_ellip)
2:Input:
3:  x_y_zx\_y\_z: Camera direction vector
4:  upup: Camera "up" pointing direction vector
5:  distdist: Distance from camera to object center
6:  FOVFOV: Camera field of view
7:  imageSizeimageSize: Desired square image dimension in pixels
8:  radrad: Characteristic radius of the small body
9:  ellip_shapeellip\_shape: Scaling factors $(a,b,c)$ matrix defining ellipsoid shape
10:  R_ellipR\_ellip: Rotation angle about the chosen rotation axis
11:  R_axisR\_axis: Rotation axis vector in body frame
12:Output:
13:  A grayscale image of the half-illuminated ellipsoid approximating the small body.
14:  Generate object (sphere, ellipsoid or mesh) $[X,Y,Z]$ coordinates using center=$(0,0,0)$ and shape=ellip_shapeellip\_shape
15:  Rotate Object by axis=R_axisR\_axis and angle=R_ellipR\_ellip
16:  Create a figure with a black background and invisible, equal, and tight axes
17:  Plot object with invisible edges and white color
18:  Set viewDirection=x_y_zx\_y\_z
[…]
20:  Set cameraUpDirection=upup
21:  Set light at infinite distance with direction=[100][-100] and Gouraud type
22:  Set dull material with 0 ambient and specular, full diffuse reflectivity and shininess coefficient of 10
23:  Set figure size = imgSizeimgSize and axis limits = [imgSize/2-imgSize/2,$imgSize/2$]
24:  Apply camera field of view = FOVFOV
25:  Convert image to grayscale
26:  Apply noise to the image
27:  Return the image
28:end procedure
A few examples of generated images are shown in Fig.˜14
Figure 14: Example synthetic images of a half-illuminated sphere with noise (left), and ellipsoid (right)
A gradient-based approach detects the horizon points from the generated synthetic images. First, calculate the gradient of the image matrix and find the gradient magnitude at each pixel. Next, find the maximum gradient magnitude across the entire image. Then, multiply this value by 0.7 to define a minimum gradient threshold and find the pixel coordinates of all points with a gradient magnitude above it. These are the desired horizon points. The value of 0.7 was chosen for the best performance after some experimental testing. To avoid detecting random noisy pixels, the absolute minimum gradient threshold was set to 0.1, which was determined experimentally by finding an appropriate threshold above the maximum gradient in dark-side lighting conditions but below the maximum gradient of regular lighting conditions. The edge detection process is expressed in Algorithm˜2:
Algorithm 2 Gradient Edge Detection Procedure
procedure pointspoints=GradientEdgeDetect($img,threshold$)
2:Input:
  imgimg: A grayscale image of the half-illuminated ellipsoid approximating the small body.
4:  thresholdthreshold: The gradient threshold used to determine edge points.
Output:
6:  pointspoints: A $3\times n$ matrix containing the coordinates of the lit edge pixels detected with each pixel stored in the form $x,y,1]^{\top}$
  Convert imgimg to numerical double format (img_dblimg\_dbl) for gradient calculation
8:  Calculate the 2D gradientgradient of img_dblimg\_dbl
  Compute the gradient_magnitudegradient\_magnitude from the gradientgradient
  Add all points with gradient_magnitude>minimum_gradientgradient\_magnitude>minimum\_gradient to pointspoints
12:  Return pointspoints
end procedure
To validate the image generation and edge detection methods, sample images of Mercury from the MESSENGER mission and corresponding distance and camera settings were used to compare results, with the OpNav measurement difference when applied on either image remaining within $1\%$ of the absolute distance. Fig.˜15 shows the similarity between the aforementioned synthetic methods and a real image.
Figure 15: Sample image comparison between PIA10172 from MESSENGER [75] of Mercury and a synthetic generated image at a distance of $\approx 27000\;\mathrm{km}$, FOV of $10.5^{\circ}$ and resolution of $1024\times 1024$, with detected horizon points overlaid.
A.1 Edge Detection Threshold Analysis
This implementation uses a pixel gradient method to detect the small body’s lit limb edge. The maximum gradient on the image is found. All pixels with a gradient within a certain threshold of the maximum gradient are stored as the detected edge points for the OpNav measurement. The selection of this gradient, a value between 0 and 1, is critical to the measurement performance. Thus, we analyzed the performance for varying thresholds. The difference between the limb detected by a high threshold and the ideal threshold is shown in Fig.˜16. Lower thresholds were found to lead to a higher lit limb angle and number of detected points, coupled with a lower analytical covariance. Using a lower threhshold may also increase the risk of misdetecting background noise or excess points (such as those along the terminator line) leading to inconsistencies between the analytical and measurement covariance. The opposite effect holds for higher thresholds. Its error did not exceed the measurement covariance as lower thresholds may have due to detecting an excess number of points or including background noise. The ideal balance was found to be a value of 0.4 as it had consistently improved performance from higher thresholds, particularly for ellipsoid cases. At this threshold, sufficient points are included to characterize the small body silhouette, while avoiding inconsistencies in analytical and measured covariances or detecting erroneous points.
[…]
Appendix B OpNav Validation with Elliptic Bodies
In this section, we validate our synthetic image generation, edge detection and OpNav implementation for consistency across spherical and elliptical bodies to support correct implementation and provide repeatability reference.
B.1 Elliptic Shape Performance
We now test the OpNav performance against increasingly elliptical bodies. The shape ratios are represented using the notation [a,b,c] where a, b and c represent the principal axis dimensions of the triaxial ellipsoid body and the principal axis frame is defined such that $a\geq b\geq c$ according to convention [16]. To isolate the effect of body ellipticity, rotation is ignored for this test. As can be seen in Fig.˜17, the OpNav error and covariance increase with highly elliptic bodies. The error remains within $\approx 3\%$ of the absolute distance and is correctly bounded by the covariance for all cases including an ellipsoid with a shape ratio of [5,1,1], supporting its usability for state estimation with ellipsoid bodies. For this test case, we simulate the spacecraft camera moving along one revolution of an FTO around the small body with a radius of 2.04292.0429km and orbit period of 3 days.
B.2 Rotating Ellipsoid Performance
The elliptic shape scenario is repeated with rotation to determine its effect on OpNav performance. A [2.5,1,1] ellipsoid with the rotation period and axis of Bennu are used as described in appendix˜A. The axis is at approximately 180 [76] from the z-axis in the Hill frame with a synodic rotation period of 4.296057 hours [68]. For this test case, we simulate along one revolution of an FTO with a radius of 2.04292.0429km. The result in Fig.˜18 suggests that rotation significantly varies the accuracy of OpNav due to various alignments leading to lit limb geometries that are better or worse for accuracy. High errors occur more frequently at higher ellipse ratios not due to limitations of the CRA, but rather due to camera limitations, as the limb reaches the limit of what can be accurately characterized with the chosen camera settings, particularly the FOV and resolution, causing relatively poorer curvature resolution or the limb exceeding the image frame. The error and covariance vary periodically with the rotation which corresponds with the higher frequency behaviour. This occurs as the limb oscillates between eccentric and circular when viewing at a perpendicular angle. The lower frequency behaviour is an effect of viewing the orbiting body from different elevations relative to the rotation axis. When straight-on to the pole, the limb curvature and error remain consistent, which corresponds to the dips in error while viewing at a perpendicular angle results in the general peaks and greatest oscillation. The measurement error was found to be correctly bounded in either rotating or non-rotating case but with increased magnitude for the worst error, reaching $\approx 10\%$ or higher levels than the non-rotating [5,1,1] ellipsoid test case. However, since the error is still correctly bounded by the covariance, the measurement can be used in rotating cases and accounted for through filtering methods by using the covariance knowledge.
Appendix C Path Constraint Partial Derivatives
@@
This section extends the derivation in Section˜3.1 by computing the partial derivatives of the path constraints. The path constraint partial derivatives in terms of Milankovitch orbital elements are as follows:
@@
Appendix D Controller Gain Tuning Guidelines
Since the gains determined for the controller used in this research are for specific test cases with Bennu’s properties, it is imperative to develop guidelines to derive it for different scenarios to make it useful for diverse missions. After investigating the relations between the gains and the controller performance, the following observations and guidelines are made:
@@ Appendix D Controller Gain Tuning Guidelines
This Lyapunov controller formulation may be adapted for different use cases using these observations. Results demonstrating the effectiveness of these gains can be seen in Section˜5.
@@
Funding
On behalf of all authors, the corresponding author states that there was no source of funding for this work.
Data Availability
Data sets generated during this study during simulations and that support its findings are available from the corresponding author on reasonable request.
Author Contributions
Both authors jointly contributed to the conceptualization and development of the ideas, theoretical framework and derivations. Aditya Arjun Anibha led the implementation, conducted the simulations, and drafted the manuscript. Dr. Kenshiro Oguri supervised the analysis, provided critical feedback, and contributed to the revision and refinement of the manuscript.